A random sample of size 80 is taken from a population that follows the following density function: x e 1280 Using the Central Limit Theorem, find the approximate probability that the sample mean exceeds 13.
Answer #1

for above is a gamma distribution for which parameter \alpha=3 and \beta=4

therefore mean =\alpha\beta =12

ad std deviation =sqrt(\alpha\beta2) =sqrt(48)=6.928

for sample size n=80 ;

estimated mean =12

and std error of mean =6.928/sqrt(80)=0.775

hence from central limit theorum ; approximate probability =P(Xbar>13)=1-P(Xbar<13)

=1-P(Z<(13-12)/0.7746)=1-P(Z<1.29)=1-0.9015=0.0985